OpenStudy (anonymous):

Convert (-5,pi/4) to rectangle coordinates. a. (5sqrt2/2,5sqrt2/2) b. (sqrt2/2,sqrt2/2) c. (5sqrt3/2,5/2) d. (-5sqrt2/2,-5sqrt2/2)

OpenStudy (ipwnbunnies):

\[Polar coordinates are \in (r,\theta)\] \[x = r*\cos \theta\] \[y = r * \sin \theta\]

OpenStudy (ipwnbunnies):

The coordinates will be written as usual: (X,Y)

OpenStudy (anonymous):

\[x=-5\cos \pi/4 \]\[y=-5\sin \pi/4\]

OpenStudy (ipwnbunnies):

Yep. Do you know both the cosine and sine of pi/4? They're the same.

OpenStudy (anonymous):

\[\sqrt{2}/2\]

OpenStudy (ipwnbunnies):

There ya go. You got this problem by the balls.

OpenStudy (anonymous):

lol so b is the answer?

OpenStudy (anonymous):

can you help me with one more question?

OpenStudy (anonymous):

or try to

OpenStudy (ipwnbunnies):

No, silly! You forgot to multiply by -5 for both x and y.

OpenStudy (ipwnbunnies):

x = -5*sqrt(2)/2 y = -5*sqrt(2)/2

OpenStudy (anonymous):

oh crap. ok

OpenStudy (ipwnbunnies):

I can try to help.

OpenStudy (anonymous):

Find the product of the complex numbers. Express your answer in trigonometric form. z1 = 7(cos(15) + isin(15)) z2 = 2(cos(110) + isin(110)) z1 = 14(cos(125) + isin(125)) z1 = 9(cos(125) + isin(125)) z1 = 5(cos(-95) + isin(-95)) z1 = 14(cos(265) + isin(265))

OpenStudy (ipwnbunnies):

Ooooo, not help with this. Not an irrational number guy.

OpenStudy (ipwnbunnies):

Sorry.

OpenStudy (anonymous):

lol its ok. thank you though!!!

OpenStudy (ipwnbunnies):

Unreal number guy I mean lol